# When do World Models Successfully Learn Dynamical Systems?

### v1

@@ Abstract
In this work, we explore the use of compact latent representations with learned time dynamics (‘World Models’) to simulate physical systems. Drawing on concepts from control theory, we propose a theoretical framework that explains why projecting time slices into a low-dimensional space and then concatenating to form a history (‘Tokenization’) is so effective at learning physics datasets, and characterise when exactly the underlying dynamics admit a reconstruction mapping from the history of previous tokenized frames to the next. To validate these claims, we develop a sequence of models with increasing complexity, starting with least-squares regression and progressing through simple linear layers, shallow adversarial learners, and ultimately full-scale generative adversarial networks (GANs). We evaluate these models on a variety of datasets, including modified forms of the heat and wave equations, the chaotic regime 2D Kuramoto–Sivashinsky equation, and a challenging computational fluid dynamics (CFD) dataset of a 2D Kármán vortex street around a fixed cylinder, where our model is successfully able to recreate the flow.
@@ 1 Introduction
World models were introduced by Ha and Schmidhuber [1] as dynamical generative models based on a latent representation of the world. Typically, a highly complex state of the environment is encoded in a sequence of latent variables by some encoder or tokenization map. LeCun explicitly demands an accurate representation of the world’s physical laws by the world model output [2]. Motivated from the success of large language models, the latent state dynamics is then modelled in an autoregressive fashion, where the next latent state is autoregressively predicted based on a history of prior latent states. Ultimately, the full state is reconstructed by some super-resolution module, which takes a sequence of latent variables as input. In this way, world models define the present state-of-the-art in the generation of dynamic scenes. Computer vision applications of word models have produced stunning success in video generation [3, 4].
Neural operator learning, on the other hand, aims for an efficient generation of solutions to physical systems governed by partial differential equations [5, 6, 7, 8]. Neural operators are trained on sample data produced by numerical simulation and then generalise to new initial conditions, from which solutions to the physical equations are generated.
Generative modelling of physical systems has been widely explored with an emphasis on turbulent flows, see e.g. [9, 10]. A few works, however, have already explored the spatio-temporal generative modelling as presented in Section 2. Here, the majority of works rely on generative learning methods that are additionally provided with physical input parameters. In particular, diffusion models trained in an autoregressive manner are purely data-driven but come with cost of slower, multi-step inference.
@@
    We introduce a system-theoretic framework to understand when world models can reliably learn dynamics from tokenized observations. This is demonstrated through experiments on a hierarchy of dynamical systems, starting from the heat and wave equations to complex turbulent flow synthesis. The study is mathematically grounded, explaining operator learning within world models, and proving that neural networks struggle to capture the time-evolution operator in systems that lack observability.
@@
    To the best of our knowledge we are the first using a purely data-driven GAN-based approach for spatial-temporal generative modelling.
@@
    We provide numerical studies on the length of token history in autoregressive predictions and super-resolution reconstruction, which is a first study of autoregressive model order reduction.
@@
    We show that world-model reconstructions can achieve low error predictions in capturing the key temporal correlations in the case of turbulence modelling, at significantly less computational cost than traditional methods.
Outline.
The remainder of the paper is structured as follows. Section 2 gives a brief review on key studies related to this work. In Section 3 we paint the theoretical backdrop to this paper and present our theoretical results. Section 4 introduces the datasets used in our experiments. Section 5 provides empirical validation of the theoretical findings, evaluating model performance on the datasets. We conclude in Section 6 and address some limitations of this work. A significant amount of extraneous material is included in the Appendix A.
Video visualisations of our solutions can be found in the supplementary material.
@@ 2 Related Work
Super-resolution reconstruction of turbulent flow via generative learning.
A prominent application area for generative learning in turbulence is resolution reconstruction, in which GANs have become the dominant approach. Numerous studies have demonstrated GANs ability to synthetically reconstruct high-resolution turbulent flow fields from low-resolution or noisy inputs [11, 12, 13, 14, 15, 16]. Most of these studies use the (Enhanced) Super-Resolution GAN ((E)SRGAN) [17, 18] incorporating physics-based loss functions to better align the reconstructions with physical laws [19, 20, 21, 22]. Unlike our proposed methods, these (E)SRGAN-based frameworks integrate physical constraints through tailored loss modifications. Similarly, approaches targeting the reconstruction of high-fidelity turbulence from low-fidelity or degraded data utilise physics-informed variants of the (E)SRGAN framework [23, 24, 25, 26]. However, GANs are not the only method for super-resolution reconstruction. Zhang et al. [8] recently introduced a novel approach based on operator learning to reconstruct high-resolution turbulent fields from sparse data using a tokenizer and energy transformer. A common limitation of these methods is that they operate on individual snapshots of the flow, meaning they cannot capture or model temporal dynamics.
Spatio-temporal generative modeling for turbulent flow.
Recent studies have addressed the issue of modelling spatio-temporal turbulent flows using generative frameworks, employing various strategies. One approach involves Transformer-based frameworks that use a Tokenizer to encode interactions between flow states over time into a latent space representation, which can be used to model turbulent flows. To improve the quality of the synthesised turbulent flows and strengthen the robustness of network training, these frameworks also incorporate Navier–Stokes constraints [27] or the spatial locations of the solution states [28, 29, 30, 31]. Another prominent framework in this research area is the recurrent neural network (RNN), which facilitates data-driven turbulence modelling by incorporating physical knowledge into the framework, for example by hard-coding known physical conditions as boundary conditions [32, 33, 34, 35] or by using pretrained networks [36]. Guidance on physical constraints also produced successful results when capturing flow dynamics using residual neural networks [37] or diffusion models [38]. The latter are particularly useful for autoregressive flow state prediction, where physics-informed learning can also be incorporated [39]. Unlike previous methods, however, diffusion models can also be used in a purely data-driven manner [40, 41]. This is likely due to the iterative nature of diffusion models, which comes at the cost of slower, multi-step inference compared to deterministic, single-pass models. McCabe et al. [42] introduced another autoregressive pretraining strategy for developing multi-use physical surrogates, which is limited to uniform grids but also does not require the incorporation of any physical parameters. Lastly, Afzaeli et al. [43] demonstrated that a combination of GANs and Autoencoders (AE) can predict fluid flow at the subsequent timestep of a given input state.
@@ 3 Theoretical Results
3.1 Dynamical systems, tokenization and autoregressive dynamics in latent space
@@
where the state space $\mathcal{X}$ is compact with differentiable boundary and $f\colon\mathbb{R}^{n}\to\mathbb{R}^{n}$ is Lipschitz with constant $L>0$ and fulfils $f(x)\cdot\nu(x)=0$ for $x\in\partial\mathcal{X}$, where $\partial\mathcal{X}$ is the boundary of $\mathcal{X}$ and $\nu(x)$ the outward normal vector field. The Picard–Lindelöf theorem [44] ensures global existence and uniqueness of solutions. Furthermore, as $f(x)$ is tangential to $\mathcal{X}$ at the boundary, the solution never leaves $\mathcal{X}$. If $x(t)$ and $\tilde{x}(t)$ start from $x_{0}$ and $\tilde{x}_{0}$, the Grönwall estimate implies continuous dependence on initial data.
Let $h\colon\mathcal{X}\to\mathbb{R}^{m}$ ($m\leq n$) be a continuous tokenization map which maps the high dimensional state $x\in\mathcal{X}$ into the latent representation $y=h(x)\in\mathbb{R}^{m}$ with $m\ll n$. In system theory $h$ is known as the (system) output map. We define the tokenized dynamics
@@
Denote by $S_{\Delta}\colon\mathcal{X}\to\mathcal{X}$ the time–$\Delta$ flow (“propagator”) associated with (1):
@@
Observability concepts for nonlinear control systems have been rigorously analysed in, e.g., [45]. Further details on observability in the standard, continuous setting of control theory can be found in Appendix A.2.
In the context of world models, the reconstruction map $G$ is also called a decoder or super-resolution map. Observability (i) and (ii) are closely related, as the following Lemma shows:
@@
Figure 1: The full-state time-series data $x_{t-k:t}=(x(t-k),\ldots,x(t))$ is tokenized via $y_{\mathrm{seq}}(t)=h(x_{t-k:t})$, yielding a low-dimensional representation. The next observation is predicted by $y(t+1)=g(y_{\mathrm{seq}}(t))$, then appended—dropping the oldest entry— to form $y_{\mathrm{seq}}(t+1)$, and finally the full state is reconstructed via $x(t+1)=G(y_{\mathrm{seq}}(t+1))$.
@@
A linear dynamical system is given by $S_{\Delta}(x)=e^{\Delta A}x$ where $A\in\mathbb{R}^{n\times n}$ is the generator of the dynamics, i.e. in (1) we have $f(x)=Ax$. In this easy case, we can drop the compactness condition for the state space $\mathcal{X}$ and identify it with $\mathbb{R}^{n}$. If we additionally have a linear observation $h\colon\mathbb{R}^{n}\to\mathbb{R}^{m},h(x)=hx$ with $h\in\mathbb{R}^{m\times n}$, observability is well-known to be equivalently characterized in terms of algebraic rank conditions [46, Chapter 6]. In particular, the seminal work of Kalman [47] related observability to the (Kalman) observability matrix which for $k\in\mathbb{N}$ is defined as
@@
For the proof of this standard result see, e.g., [46, Chapter 3]. Note that the third condition is typically referred to as the (Popov-Belevitch-)Hautus test for observability. ∎
Unfortunately, the situation in the non-linear case is less well understood. We refer to Appendix A.2 for a detailed discussion of the state of research.
@@ 3.2 Operator learning using the world model approach
where $\eta_{g}:\mathcal{Y}^{\times k}\to\mathcal{Y}$ and $\eta_{G}:\mathcal{Y}^{\times k}\to\mathcal{X}$ are represented by neural networks from hypothesis spaces $\mathcal{H}_{g,N,T}$ and $\mathcal{H}_{G,N,T}$, respectively and $U(A)$ stands for the uniform distribution on the finite set $A$. Let ^$\hat{\mathcal{L}}_{N,T,g}(\eta_{g})$ and ^$\hat{\mathcal{L}}_{N,T,G}(\eta_{G})$ be the corresponding empirical risk function learning from $N$ trajectories of length $T$, see Appendix A.1 for the definition. For the neural networks $\hat{\eta}_{g,N,T}$ and $\hat{\eta}_{G,N,T}$ trained by empirical risk minimization
we obtain the following theorem on probably approximately correct (PAC) learning, see Appendix A.1 for the proofs of this section.
@@
$P(\mathcal{L}_{g,N(\varepsilon,\delta),T}(\hat{\eta}_{g,N(\varepsilon,\delta),T})\leq\varepsilon)\geq 1-\delta\text{ and }P(\mathcal{L}_{G,N(\varepsilon,\delta),T}(\hat{\eta}_{G,N(\varepsilon,\delta),T})\leq\varepsilon)\geq 1-\delta$ (9)
Let us next consider a random state $x_{0}\sim\mu$ and its past token sequence $y_{\mathrm{past}}(x_{0})$ of length $k$ and $y_{\mathrm{past}-}(x_{0})$ on length $(k-1)$. We define the risk function measuring the error in the autoregressive update of the full state $S_{\Delta}(X)$ as follows
@@
In the case that (1) defines an ergodic dynamical system with invariant measure $\mu$, we can replace the large sample limit in the number of trajectories $N(\varepsilon,\delta)\to\infty$ as $\varepsilon,\delta\downarrow 0$ with the large time limit $T(\varepsilon,\delta)\to\infty$ keeping $N=1$ fixed, see [10] for this approach in a related situation based on Birkhoff’s ergodic theorem. The autoregressive approach can also be trained variationally in GAN-fashion, see Appendix A.3 for the details.
@@ 3.3 Examples
The examples we consider are spatially or spatio-temporally discretized partial differential equations (PDE) on a quadratic lattice, see Appendix A.4. In all cases we utilize a simple tokenizer used e.g. in [48], taking the average on patches of pixel size $\sqrt{\kappa}\times\sqrt{\kappa}$ with $\sqrt{\kappa}\in\mathbb{N}$ and $m=n/\kappa\in\mathbb{N}$.
@@
Our experiments in Fig. 2 show a significant deterioration of operator learning with world models for the case of the heat- and wave equation with constant coefficients. To produce positive results, we thus perturb the lattice Laplacian using a non constant eigenfunction $a(\xi)=\exp(Z(\xi))$, where $Z(\xi)$ is a restriction of a realization of a Gaussian random field to $\Gamma$. Numerical tests based on diagonalization of the so-obtained matrix show that generally the eigenvectors are not annihilated by the tokenizer $h$ for non constant coefficients. Correspondingly, numerical errors are much reduced.
@@
[…]
@@
Following [10][Eq. (24)], in this experiment, we use a non-linear tokenization which only observes the turbulence strength, which measures the deviation of the absolute velocity from the long time average. This gives a scalar turbulence field, to which thereafter the standard tokenizer of LongVideoGAN [48] is applied. The reconstruction map $G$ is not considered for the full state, but for the full state projected to the turbulence strength, which is a continuous mapping. Observability of this projected state follows from the observability of the full state, which is unknown. The quality of the world model reconstructions thus has to serve as experimental evidence of observability.

### v2

@@ Abstract
In this work, we explore the use of compact latent representations with learned time dynamics (‘World Models’) to simulate physical systems. Drawing on concepts from control theory, we propose a theoretical framework that explains why projecting time slices into a low-dimensional space and then concatenating to form a history (‘Tokenization’) is effective at learning physics datasets, and characterise when exactly the underlying dynamics admit a reconstruction mapping from the history of previous tokenized frames to the next. To validate these claims, we develop a sequence of models with increasing complexity, starting with least-squares regression and progressing through simple linear layers, shallow adversarial learners, and ultimately full-scale generative adversarial networks (GANs). We evaluate these models on a variety of datasets, including modified forms of the heat and wave equations, the chaotic regime 2D Kuramoto-Sivashinsky equation, and a challenging computational fluid dynamics (CFD) dataset of a 2D Kármán vortex street around a fixed cylinder, where our model is successfully able to recreate the flow. Comparisons to FNO and DeepONet show comparable short term and improved long term accuracy of solutions generated by world models.
@@ 1 Introduction
World models were introduced by ha2018world as dynamical generative models based on a latent representation of the world. Typically, a highly complex state of the environment is encoded in a sequence of latent variables by some encoder or tokenization map. lecun2022path explicitly demands an accurate representation of the world’s physical laws by the world model output. Motivated from the success of large language models, the latent state dynamics are then modelled in an autoregressive fashion, where the next latent state is autoregressively predicted based on a history of prior latent states. Ultimately, the full state is reconstructed by some super-resolution module, which takes a sequence of latent variables as input. In this way, world models define the present state-of-the-art in the generation of dynamic scenes. Computer vision applications of word models have produced stunning success in video generation (hu2023gaia; peng2025open).
Neural operator learning, on the other hand, aims for an efficient generation of solutions to physical systems governed by partial differential equations (li2020fourier; lu2021learning; kovachki2023neural; zhang2025operator). Neural operators are trained on sample data produced by numerical simulation and then generalise to new initial conditions, from which solutions to the physical equations are generated.
Generative modelling of physical systems has been widely explored with an emphasis on turbulent flows, see e.g. Kim:2020; drygala2022generative.
World models have been applied previously to simulate physical systems (skorokhodov2022stylegan; klemmer2022spate), but their focus is on video generation, so they neither analyse system-theoretic foundations nor compare with operator-learning approaches.
@@
    We introduce a system-theoretic framework to understand when world models can reliably learn dynamics from tokenized observations. This is demonstrated through experiments on a hierarchy of dynamical systems, starting from the heat and wave equations to complex turbulent flow synthesis. The study is mathematically grounded, explaining operator learning within world models, and proving that neural networks struggle to capture the time-evolution operator in systems that lack observability. Video visualisations of our solutions can be found in the supplementary material.
@@
    We interpret world models as an operator learning method and compare this approach to state of the art methods like FNO and DeepONet showing superior long time stability of world models for in-distribution and out-of-distribution test data.
@@
    We provide numerical studies on the length of token history in autoregressive predictions and super-resolution reconstruction, which is a first study of autoregressive model order reduction for dynamical systems.
@@
    We show that world-model based simulation can achieve low error predictions in capturing the key temporal correlations in the case of turbulence modelling even in situations where the state cannot be fully observed, at significantly less computational cost than traditional methods from computational fluid dynamics.
Section 2 discusses related work, Section 3 the theoretical setting, Section 4 datasets and experiments, Section 5 our results, which we discuss in Section 6. We will publish our code after acceptance.
@@ 2 Related Work
Generative super-resolution reconstruction of turbulent flow
A prominent application of generative learning in turbulence is resolution reconstruction. GAN-based methods dominate this area, with (Enhanced) Super-Resolution GAN (wang2018esrgan) variants widely used to recover high-resolution turbulent fields from low-resolution or noisy inputs, often enhanced by physics-informed loss functions (deng2019super; bode2023applying; chen2024reconstructing; zheng2024high). More recently, (zhang2025operator) proposed an operator-learning approach that reconstructs high-resolution fields from sparse data using a tokenizer and energy transformer.
Spatio-temporal generative modelling for turbulent flow.
Moving beyond snapshot reconstruction, recent studies have applied generative learning to spatio-temporal turbulence modelling. Transformer-based frameworks encode temporal interactions into latent spaces, guided by Navier–Stokes constraints (xu2025amr) or spatial embeddings (li2022transformer; alkin2024universal). Recurrent neural networks integrate physics through boundary conditions (ren2023physr; ren2025learning) or pretrained modules (wan2025pimrl). Residual networks (liu2024multi) and diffusion models (NEURIPS2023_8df90a14) have also been applied, with diffusion models notable for operating in a purely data-driven manner (kohl2024benchmarking; yang2023diffusion), but limited by slower inference. Other directions include autoregressive pretraining for surrogates (mccabe2024multiple), GAN–Autoencoder hybrids for next-step prediction (afzali2021latent), and world models adapted from video generation (skorokhodov2022stylegan; klemmer2022spate).
Operator learning
focuses on approximating mappings between infinite-dimensional function spaces, allowing models to capture entire solution operators rather than just pointwise predictions. Approaches such as Physics-informed Neural Networks (PINN) (raissi2019physics) and Neural ODE (chen2018neural) utilise neural networks to solve specific ordinary or partial differential equations (ODE/PDE), either by embedding physical laws into the loss function or by parametrising continuous-time dynamics. These methods are generalised by neural operator learning, which aims to efficiently generate solutions for entire families of PDEs by instead learning solution operators, which take as input both the problem specification (e.g. coefficients, forcing terms, boundary or initial conditions) and the associated function space, and output the corresponding solution function. Within this framework, Fourier Neural Operators (FNO) (li2020fourier; kovachki2023neural) perform convolutions in the spectral domain to capture long-range interactions, while DeepONets (lu2021learning) employ a branch-trunk architecture to flexibly map input functions to outputs. For a comprehensive overview of (neural) operator learning methods and frameworks, see cai2021physics; tanyu2023deep; kovachki2024operator
@@ 3 Theoretical Results
3.1 Dynamical systems, tokenization and autoregressive dynamics
@@
where the state space $\mathcal{X}$ is compact with differentiable boundary and $f\colon\mathbb{R}^{n}\to\mathbb{R}^{n}$ is Lipschitz with constant $L>0$ and fulfils $f(x)\cdot\nu(x)=0$ for $x\in\partial\mathcal{X}$, where $\partial\mathcal{X}$ is the boundary of $\mathcal{X}$ and $\nu(x)$ the outward normal vector field. The Picard–Lindelöf theorem (picard) ensures global existence and uniqueness of solutions. Furthermore, as $f(x)$ is tangential to $\mathcal{X}$ at the boundary, the solution never leaves $\mathcal{X}$. If $x(t)$ and $\tilde{x}(t)$ start from $x_{0}$ and $\tilde{x}_{0}$, the Grönwall estimate implies continuous dependence on initial data.
Let $h\colon\mathcal{X}\to\mathbb{R}^{m}$ ($m\leq n$) be a continuous tokenization map which maps the high dimensional state $x\in\mathcal{X}$ into the latent representation $y=h(x)\in\mathbb{R}^{m}$ with $m\ll n$. In system theory $h$ is known as the (system) output map. We define the tokenized (latent) dynamics
@@
[…]
@@
Observability concepts for nonlinear control systems have been rigorously analysed in, e.g., Controllability_and_Observability. Further details on observability in the standard, continuous setting of control theory can be found in Appendix B.
In the context of world models, the reconstruction map $G$ is also called a decoder or super-resolution map (Fig. 1). Observability (i) and (ii) are closely related, as the following Lemma shows:
@@
Figure 1: The full‐state time‐series data $x_{t-k:t}=(x(t-k),\ldots,x(t))$ is tokenized via $y_{\mathrm{seq}}(t)=h(x_{t-k:t})$, yielding a low‐dimensional representation. The next observation is predicted by $y(t+1)=g(y_{\mathrm{seq}}(t))$, then appended—dropping the oldest entry— to form $y_{\mathrm{seq}}(t+1)$, and finally the full state is reconstructed via $x(t+1)=G(y_{\mathrm{seq}}(t+1))$.
@@
A linear dynamical system is given by $S_{\Delta}(x)=e^{\Delta A}x$ where $A\in\mathbb{R}^{n\times n}$ is the generator of the dynamics, i.e. in (1) we have $f(x)=Ax$. In this easy case, we can drop the compactness condition for the state space $\mathcal{X}$ and identify it with $\mathbb{R}^{n}$. If we additionally have a linear observation $h\colon\mathbb{R}^{n}\to\mathbb{R}^{m},h(x)=hx$ with $h\in\mathbb{R}^{m\times n}$, observability is well-known to be equivalently characterized in terms of algebraic rank conditions (Son98, Chapter 6). In particular, the seminal work of Kal63 related observability to the (Kalman) observability matrix which for $k\in\mathbb{N}$ is defined as
@@
For the proof of this standard result see, e.g., (Son98, Chapter 3). Note that the third condition is typically referred to as the (Popov-Belevitch-)Hautus test for observability. ∎
Unfortunately, the situation in the non-linear case is less well understood. We refer to Appendix B for a detailed discussion of the state of research.
@@ 3.2 Operator learning using the world model approach
where $\eta_{g}:\mathcal{Y}^{\times k}\to\mathcal{Y}$ and $\eta_{G}:\mathcal{Y}^{\times k}\to\mathcal{X}$ are represented by neural networks from hypothesis spaces $\mathcal{H}_{g,N,T}$ and $\mathcal{H}_{G,N,T}$, respectively and $U(A)$ stands for the uniform distribution on the finite set $A$. Let $\hat{\mathcal{L}}_{N,T,g}(\eta_{g})$ and $\hat{\mathcal{L}}_{N,T,G}(\eta_{G})$ be the corresponding empirical risk function learning from $N$ trajectories of length $T$, and $\hat{\eta}_{g,N,T},\,\hat{\eta}_{G}$ where they take their respective minimums. See Appendix A for definitions.
We obtain the following theorem on probably approximately correct (PAC) learning, see Appendix A for the proofs in this section.
@@
[…]
Let us next consider a random state $x_{0}\sim\mu$ and its past token sequence $y_{\mathrm{past}}(x_{0})$ of length $k$ and $y_{\mathrm{past}-}(x_{0})$ on length (k1)(k-1). We define the risk function measuring the error in the autoregressive update of the full state $S_{\Delta}(X)$ as follows
@@
The autoregressive approach can also be trained variationally in GAN-fashion, see Appendix D for the details.
@@ 3.3 Examples
The examples we consider are spatially or spatio-temporally discretized partial differential equations (PDE) on a quadratic lattice, see Appendix E. In all cases we utilize a simple tokenizer map $h(x)$ used e.g. in Long_video_gan, taking the average on patches of pixel size $\sqrt{\kappa}\times\sqrt{\kappa}$ with $\sqrt{\kappa}\in\mathbb{N}$ and $m=n/\kappa\in\mathbb{N}$.
@@
See Appendix A for the proof. Our experiments in Fig. 2 show a significant deterioration of operator learning with world models for the case of the heat equation with constant coefficients. To produce positive results, we thus perturb the lattice Laplacian using a non constant eigenfunction $a(\xi)=\exp(Z(\xi))$, where $Z(\xi)$ is a restriction of a realization of a Gaussian random field to $\Gamma$. Numerical tests based on diagonalization of the so-obtained matrix show that generally the eigenvectors are not annihilated by the tokenizer $h$ for non constant coefficients. Correspondingly, numerical errors are much reduced.
@@
where $\nabla^{2}$ is the Laplace operator $\nabla\cdot\nabla$. Observability of the KSE is discussed in Appendix F.
@@
Following drygala2022generative Eq. (24), in this experiment, we use a non-linear tokenization which only observes the turbulence strength, which measures the deviation of the absolute velocity from the long time average. This gives a scalar turbulence field, to which thereafter the standard tokenizer of LongVideoGAN (Long_video_gan) is applied. The reconstruction map $G$ is not considered for the full state, but for the full state projected to the turbulence strength, which is a continuous mapping. Observability of this projected state follows from the observability of the full state, which is unknown. The quality of the world model reconstructions thus has to serve as experimental evidence of observability.